OpenStudy (anonymous):

Find the line which passes through (1,-3) and (11,2). Write your answer in slope-intercept form.

OpenStudy (anonymous):

slope(m)=(y2-y1)/(x2-x1) =2+3/11-1 =5/10=1/2 slope intercept form is given by y-y1=m(x-x1) (x1,y1)=(1,-3)..... then substituting in slope intercept form....we get y+3=1/2(x-1)...this gives the equation

OpenStudy (anonymous):

thank you.